A circular wading pool requires painting.The pool floor is 85m square. It is only 0.3m deep. How much paint would be required to paint the pool?Show work.
Found 2 solutions by nyc_function, mananth:
Answer by nyc_function(2741)

Let V = volume
V = area of base times height
V = 85 times 0.3
Can you finish?

Answer by mananth(16946)

base area of pool = 85 sq. m
area = pi^r^2
r^2= area /pi
r=sqrt(area / pi)
radius = sqrt(85 / pi)
= 5.2 m
Area of wall = 2*pi*r*h
Area of the wall 9.8
Area of wall + area of base=
5.2 + 9.8 = 15 sq. mt